OpenStudy (anonymous):

Find the ratio of AD:BD if ratio of AC:DE = 55:28

OpenStudy (anonymous):



OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

\[\frac{AC}{DE}=\frac{AB}{AD}=\frac{55}{28}=\frac{AB-AD}{AD}=\frac{55-28}{28}=\frac{DB}{AD}=\frac{27}{28}\]